Greenhab Report – December 15th

Greenhab Officer:
Walter Calles

Environmental Control:
40% Shade Cloth on

Average Temperatures:
Low: 24 C
High: 52 C

Hours of Supplemental Light:

Daily Water Usage of Crops:
6 gallons

Water in Blue Tank (lbs):

Times of Watering for Crops:

Changes to crops:

Just learned the Greenhab position and responsibilities. First day of the mission, pretty excited to grow all types of crops. Will plan for future dinner dishes.


Support/Supplies Needed:
NONE (yet!)

Sol Summary – December 15th

Crew 200 Sol Summary Report 15-DEC-2018

Sol: 13

Summary Title: Cleaning, gathering food, and two happy crews meet

Author’s name: Crew 200 and 201

Mission Status: Transitioning the habitat from Crew 200 to Crew 201.

Sol Activity Summary: One crew 200 members drove to Grand Junction this morning to pick up Crew 201! Two members from crew 200 remained in the habitat and thoroughly cleaned. Crew 200 welcomed and oriented Crew 201 and also helped them gather new food from Hanksville. A lot of delicious, dehydrated food! Both crews will enjoy dinner together tonight!

Look Ahead Plan: Crew 200 will leave for Grand Junction tomorrow and Crew 201 will eagerly begin their mission!

Anomalies in work: N/A

Weather: Chilly, snow still melting, muddy

Crew Physical Status: All crew are happy and healthy


Reports to be filed: Operations Report, GreenHab Report, Sol Report,

Support Requested: N/A

Operations Report – December 15th

SOL: 0

Name of person filing report: Cesar Serrano

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 5:00 pm last night to around 9:30 am today. On tonight at 5:00pm.

Solar— 81%
Diesel Reading – 95 %
Propane Reading – 25 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 300 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 6

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours: 141.5
Beginning charge:
Ending charge: 70%
Currently charging: no

Sojourner rover used: no
Hours: 43.1
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: no
Hours: 69.6
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: no
Hours: 43.5
Beginning charge:
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours: 58.8
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: we cleaned the mud from the rovers
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? Yes, to go to Hanksville to take water
CrewCar used and why, where? Yes, to go to Grand Junction to pick up the next crew. Refill at Hanksville (MDRS account) + at Grand Junction (Antoine account). Used then to go to Hanksville to take groceries (3rd refill, MDRS account).

General notes and comments: we cleaned the station for checkout, while Antoine went picking up the next crew in Grand Junction. Crew 200 welcomed and trained crew 201.

Summary of internet: Nothing to report

Summary of suits and radios: we repaired the spacesuit 3, according to mission support instructions.

Summary of Hab operations: Cleaning, welcome of Crew 201

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: Floor mopped and cleaned again, better.

Summary of RAM operations: Cleaned up.

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site